xsquared = negative 4

Answer by stanbon(75887)

xsquared = negative 4
---
x^2 = -4
---
x^2 + 4 = 0
(x+2i)(x-2i) = 0
x = -2i or x = 2i
====================
Cheers,
Stan H.